OpenStudy (anonymous):

What is the solution to the equation? 5+(x-12)=18

OpenStudy (anonymous):

ok so what do you think it is

OpenStudy (anonymous):

or how do you do it and work it out

OpenStudy (anonymous):

idk that's y I need help

OpenStudy (anonymous):

hang on give me a min

OpenStudy (anonymous):

ok lets do this step by step

OpenStudy (anonymous):

i will help you

OpenStudy (anonymous):

5+(x-12)=18 subtract 5 on both sides. 5-5+(x-12)=18-5 x-12=13 add 12 on both sides x-12+12=13+12 x=25

OpenStudy (anonymous):

@keepsecure93

OpenStudy (anonymous):

the answer is x= 25